OpenStudy (anonymous):

How would I solve the question below:

OpenStudy (anonymous):

For a line 10 in long, what is the probability that of 4 randomly distributed points, there will be no distance between two points less than 1 inch?

OpenStudy (anonymous):

Hmmm, interesting question. I have some theories on how to solve...

OpenStudy (anonymous):

Maybe first try to find out how you would do it with just two points.

OpenStudy (anonymous):

Then add in a third point and see how that would affect it.

OpenStudy (anonymous):

I have no clue how to solve it with just 2 points

OpenStudy (anonymous):

Okay, so think about the points and their distance from the left endpoint: 

OpenStudy (anonymous):

then we want |x_1-x_2|>1... but how do we get from that to a probability?

OpenStudy (anonymous):

We have two random variables.

OpenStudy (anonymous):

\[ \Pr(Y\geq X-1, Y\leq X+1) \]

OpenStudy (anonymous):

Hmmm, hold on for a second... I'm still actually trying to think about how to do this.

OpenStudy (anonymous):

How about we do it like this...  \[ x_1+x_2+x_3+x_4+x_5 = 10 \]